# A Third-Order Relativistic Algebraic Diagrammatic Construction Method for Double Ionization Potentials: Theory, Implementation, and Benchmark

### v1

@@ III Computational details
The relativistic DIP-ADC methods up to third order have been implemented in our in-house quantum chemistry package BAGH.Dutta et al. (2025) While the core architecture of BAGH is written in Python, computationally intensive routines are handled via Fortran and Cython for efficiency. It supports a range of non-relativistic and relativistic (4c and 2c) electronic structure methods, including many-body perturbation theory, coupled-cluster, and ADC approaches. BAGH utilizes external quantum chemistry programs for the generation of one and two-electron integrals, and presently supports interfaces with PySCF,Sun (2015); Sun et al. (2018, 2020) GAMESS-US,Barca et al. (2020) DIRAC,Jensen et al. (2022) and socutils.Wang (2025) In this work, all calculations were carried out using the PySCF interface. The X2CAMF-HF calculations were carried out using the socutils package. Cholesky vectors of LOO and LOV type are generated prior to the construction of the FNS basis and stored for subsequent use, where O and V denote occupied and virtual spinors, respectively. The LVV-type Cholesky vectors are only generated in the FNS basis. In all calculations, the Cholesky vectors were generated using a fixed decomposition threshold of $10^{-5}$. The FNS basis is then constructed from the ground-state MP2 amplitudes. Following this, LVV-type Cholesky vectors are generated directly in the FNS basis. This approach not only speeds up the integral transformation step but also greatly reduces memory usage. Finally, the DIP-ADC(3) equations in the FNS basis are solved using a Davidson iterative diagonalization scheme, with initial guess vectors obtained from DIP-ADC(1) calculations. Figure 1 provides a schematic overview of the FNS-CD-X2CAMF-DIP-ADC(3) algorithm. In the four-component and non-relativistic (NR) calculations, the CD approximation to the two-electron integrals was not employed. The dyall.avxz (x = 2, 3, 4) basis setsDyall (1998, 2002, 2006, 2016) were employed for the valence DIP calculations. Diffused versions of the dyall.av4z basis set was generated by augmenting with single and double diffuse functions, generated using the DIRACJensen et al. (2022) software package. Experimental geometries of the diatomic molecules were obtained from the NIST database.Johnson (2022) Throughout this work, the frozen core approximation was employed for all correlation calculations.
@@ IV.1 Choice of FNS threshold
The formation of doubly ionized states involves a large orbital relaxation effect. Therefore, the FNS-induced truncation of $3h1p$ configurations in DIP-ADC(3) can significantly affect the accuracy of the calculation. It is essential to benchmark the effect of the truncation on the calculated DIP values. Figure 2 shows the absolute error in DIP values, as a function of the percentage of virtual orbitals retained in the correlation treatment. The untruncated X2CAMF Hamiltonian-based DIP-ADC(3) canonical values are used as the reference. The data correspond to the lowest vertical DIP of the Cl2 molecule computed with the dyall.av4z basis set. The results clearly show that the FNS approach converges much more rapidly than the strategy based on truncating canonical virtual spinors. Even with only 40% of the frozen virtual spinors retained, the FNS-DIP-ADC(3) method yields an absolute error of less than 0.05 eV, whereas the same level of truncation in the canonical basis results in a much larger error of approximately 0.2 eV. Convergence is achieved with just 50% of the virtual spinors in the FNS basis, compared to the 70% required in the canonical case. A more convenient and systematic way to truncate the FNS basis is to apply an occupation number threshold, referred to here as the FNS threshold, which excludes all spinors with occupation numbers below a specified value. In Figure 3, we present the absolute error in the DIP energy, computed using the FNS-DIP-ADC(3) method, relative to its canonical counterpart. The error is plotted as a function of the negative logarithm of the FNS threshold for the ground state of the doubly ionized Cl2 molecule, using the dyall.av4z basis set. It can be observed that the FNS-DIP-ADC(3) results converge closely to the canonical values at a threshold of $10^{-4.5}$. Therefore, all subsequent calculations are performed using this threshold.
@@ IV.2 Benchmarking FNS-CD-X2CAMF results against 4c-DIP-ADC(3) method
To evaluate the performance of the FNS-CD-X2CAMF-based method against the four-component FNS approach, we computed vertical DIPs for Kr, Xe, Cl2, and HBr using the DIP-ADC(3) method in both frameworks with the dyall.av4z basis set. The results are presented in Table 1. The results clearly show that the FNS-CD-X2CAMF-DIP-ADC(3) method yields values nearly identical to the four-component Dirac-Coulomb results across all systems and states considered, with error not exceeding 0.001 eV. This confirms that the former approach can reproduce fully relativistic results with negligible loss in accuracy, while offering substantial gains in computational speed and memory efficiency. Accordingly, all subsequent DIP calculations in this work employ the FNS-CD-X2CAMF-DIP-ADC methods unless stated otherwise.
@@ IV.3 Choice of basis set
To understand the effect of basis sets on the calculated DIP values, we utilize the dyall.avxz (x = 2, 3, 4) series,Dyall (1998, 2002, 2006, 2016) which provides a systematically improvable basis set hierarchy. Additionally, to examine the role of diffuse functions, we consider both singly and doubly augmented versions of the dyall.av4z basis set. The results, presented in Table 2, illustrate the dependency of DIP values in different basis sets for four different doubly ionized states of the Cl2 molecule. It is evident that increasing the basis set quality from double-zeta to triple-zeta leads to an average increase of approximately 0.2 eV in the DIP values for the four considered states. A further improvement from triple-zeta to quadruple-zeta results in a slightly smaller average shift of 0.16 eV. Unfortunately, quintuple-zeta basis sets could not be tested as they are not yet available for Cl.Reitsma et al. (2025) Interestingly, the addition of diffuse functions appears to have minimal impact: the singly augmented dyall.av4z basis set yields no noticeable change in DIP values, while the doubly augmented version results in a modest increase of only 0.01 eV per state. Thus, all further computations employ the dyall.av4z basis set without additional diffuse functions.
@@ IV.4 Comparison of DIP-ADC results with available theoretical and experimental data
To benchmark the performance of DIP-ADC(2), DIP-ADC(2)-x, and DIP-ADC(3) methods, we have computed the DIPs of Ar, Kr, Xe, and Rn, for which corresponding experimental dataKramida et al. (2024) are available. The results are presented in Table 3, alongside FNS-DIP-EOM-CCSD($3h1p$) values and various statistical error metrics, with experimental data taken as the reference. For simplicity, we will hereafter denote DIP-EOM-CCSD($3h1p$) as CCSD($3h1p$) throughout this work. The statistical parameters reported include the mean absolute error (MAE), maximum absolute deviation (MAD), standard deviation (STD), and root mean square deviation (RMSD). It can be clearly seen that DIP-ADC(2) consistently underestimates the double ionization potentials for all the noble gas states studied, with a MAE of 2.17 eV. The largest deviation, 2.77 eV, is observed for the $^{1}D_{2}$ state of the Ar atom. This poor performance is not unexpected, as double ionization leads to a significant deficiency in the reference ground state, requiring substantial orbital relaxation to describe the doubly ionized states accurately. A more precise treatment of the $3h1p$ excitations, achieved by incorporating higher-order terms in the ADC secular matrix, is essential for accurately capturing these effects. Since ADC(2) only includes zeroth-order terms in the 3h1p3h1p-$3h1p$ block, it fails to account for the necessary relaxation, resulting in large errors. The first level of improvement is seen with the ADC(2)-x method, which includes first-order corrections in the 3h1p3h1p-$3h1p$ block in an ad hoc manner. This lowers the MAE from 2.17 eV, observed in strict ADC(2), to 0.58 eV in ADC(2)-x (see Table S3 of the Supplementary Material). Among all the methods tested, DIP-ADC(3) shows the best agreement with experiment, with an MAE of only 0.1 eV and a MAD of 0.43 eV. In comparison, the FNS-CCSD($3h1p$) method yields a MAE of 0.17 eV and a MAD of 0.35 eV.
Table 4 presents the doubly ionized states of linear diatomic molecules-Cl2, Br2, HBr, and HI. For the ADC(2) method, a large MAE of 1.73 eV and a MAD of 2.02 eV are observed. In contrast, the significantly lower MAE of 0.30 eV obtained from the FNS-CCSD(3h1p) method highlights the importance of incorporating higher-order treatment of $3h1p$ excitations. This is further supported by the DIP-ADC(3) results, which demonstrate excellent performance with a MAE of just 0.11 eV and an RMSD of 0.14 eV across the 15 states examined in this study. The DIP-ADC(2)-x method yields a significantly lower MAE of 0.55 eV compared to the 1.73 eV observed with the strict ADC(2) (See Table S4 of the Supplementary Material).
@@
IV.5 The level of treatment of relativistic Hamiltonians

### v2

@@ III Computational details
The relativistic DIP-ADC methods up to third order have been implemented in our in-house quantum chemistry package BAGH.Dutta et al. (2025) While the core architecture of BAGH is written in Python, computationally intensive routines are handled via Fortran and Cython for efficiency. It supports a range of non-relativistic and relativistic (4c and 2c) electronic structure
Figure 1: Schematic representation of the FNS-CD-X2CAMF-DIP-ADC(3) algorithm.
methods, including many-body perturbation theory, coupled-cluster, and ADC approaches. BAGH utilizes external quantum chemistry programs for the generation of one and two-electron integrals, and presently supports interfaces with PySCF,Sun (2015); Sun et al. (2018, 2020) GAMESS-US,Barca et al. (2020) DIRAC,Jensen et al. (2022) and socutils.Wang (2025) In this work, all calculations were carried out using the PySCF interface. The X2CAMF-HF calculations were carried out using the socutils package. Cholesky vectors of LOO and LOV type are generated prior to the construction of the FNS basis and stored for subsequent use, where O and V denote occupied and virtual spinors, respectively. The LVV-type Cholesky vectors are only generated in the FNS basis. In all calculations, the Cholesky vectors were generated using a fixed decomposition threshold of $10^{-5}$, which was found to give similar accuracy as that of the standard integrals.Chamoli et al. (2025); Chamoli, Nayak, and Dutta (2025) The FNS basis is then constructed from the ground-state MP2 amplitudes. Following this, LVV-type Cholesky vectors are generated directly in the FNS basis. This approach not only speeds up the integral transformation step but also greatly reduces memory usage. Finally, the DIP-ADC(3) equations in the FNS basis are solved using a Davidson iterative diagonalization schemeDavidson (1975), with initial guess vectors obtained from DIP-ADC(1) calculations. Figure 1 provides a schematic overview of the FNS-CD-X2CAMF-DIP-ADC(3) algorithm. In the four-component and non-relativistic (NR) calculations, the CD approximation to the two-electron integrals was not employed. The dyall.avxz (x = 2, 3, 4) basis setsDyall (1998, 2002, 2006, 2016) were employed for the valence DIP calculations. Diffused versions of the dyall.av4z basis set were generated by augmenting with single and double sets of diffuse functions, generated using the DIRACJensen et al. (2022) software package. Experimental geometries of the diatomic molecules were obtained from the NIST database.Johnson (2022) Throughout this work, the frozen core approximation was employed for all correlation calculations.
@@ IV.1 Choice of FNS threshold
The formation of doubly ionized states involves a large orbital
Figure 2: Convergence of the absolute error (in eV) in the lowest DIP value with respect to the size of the virtual space in the canonical and FNS versions of CD-X2CAMF-DIP-ADC(3) method for the Cl2 molecule. The full basis canonical CD-X2CAMF-DIP-ADC(3) result is used as reference, and the dyall.av4z basis set has been used for the calculations.
relaxation effect. Therefore, the FNS-induced truncation of $3h1p$ configurations in DIP-ADC(3) can significantly affect the accuracy of the calculation. It is essential to
Figure 3: Convergence of the absolute error in the DIP value (in eV) computed using FNS-CD-X2CAMF-DIP-ADC(3) method, with respect to the FNS threshold for Cl2 molecule. The dyall.av4z basis set has been used for the calculations, and the untruncated canonical value is taken as the reference.
Table 2: Basis set convergence of FNS-CD-X2CAMF-DIP-ADC(3) values (in eV) of Cl2 molecule in different Dyall basis sets.
State dyall.av2z dyall.av3z dyall.av4z s-aug-dyall.av4z d-aug-dyall.av4z Expt.McConkey et al. (1994)
$X^{3}\Sigma^{-}$ 30.90 31.11 31.27 31.27 31.28 31.13
$a^{1}\Delta$ 31.44 31.62 31.77 31.77 31.78 31.74
$b^{1}\Sigma^{+}$ 31.80 32.00 32.16 32.16 32.17 32.12
$c^{1}\Sigma^{-}$ 32.82 33.01 33.17 33.17 33.18 32.97
benchmark the effect of the truncation on the calculated DIP values. Figure 2 shows the absolute error in DIP values, as a function of the percentage of virtual orbitals retained in the correlation treatment. The untruncated X2CAMF Hamiltonian-based DIP-ADC(3) canonical value is used as the reference. The data correspond to the lowest vertical DIP of the Cl2 molecule computed with the dyall.av4z basis set. The results clearly show that the FNS approach converges much more rapidly than that obtained by truncating canonical virtual spinors. Even with only 40% of the frozen natural virtual spinors retained, the FNS-DIP-ADC(3) method yields an absolute error of less than 0.05 eV, whereas the same level of truncation in the canonical basis results in a much larger error of approximately 0.2 eV. Convergence is achieved with just 50% of the virtual spinors in the FNS basis, compared to the 70% required in the canonical case. A more convenient and systematic way to truncate the FNS basis is to apply an occupation number threshold, referred to here as the FNS threshold, which excludes all spinors with occupation numbers below a specified value. In Figure 3, we present the absolute error in the DIP energy, computed using the FNS-DIP-ADC(3) method, relative to its canonical counterpart. The error is plotted as a function of the negative logarithm of the FNS threshold for the ground state of the doubly ionized Cl2 molecule, using the dyall.av4z basis set. It can be observed that at a threshold of $10^{-4.5}$, the FNS-DIP-ADC(3) result almost converges to the canonical value. Therefore, all subsequent calculations are performed using this threshold.
@@ IV.2 Benchmarking FNS-CD-X2CAMF results against 4c-DIP-ADC(3) method
To evaluate the performance of the FNS-CD-X2CAMF-based method against the four-component FNS approach, we computed vertical DIPs for Kr, Xe, Cl2, and HBr using the DIP-ADC(3) method in both frameworks with the dyall.av4z basis set. The results are presented in Table 1. The results clearly show that the FNS-CD-X2CAMF-DIP-ADC(3) method yields values nearly identical to the four-component Dirac-Coulomb results across all systems and states considered, with error not exceeding 0.001 eV. This confirms that the former approach can reproduce four-component relativistic results with negligible error, while offering substantial gains in computational speed and memory efficiency. Consequently, all subsequent DIP calculations in this work employ the FNS-CD-X2CAMF-DIP-ADC methods.
@@ IV.3 Choice of basis set
Table 3: Comparison of error in DIP values (in eV) using 4c-FNS-CCSD($3h1p$), FNS-CD-X2CAMF-ADC(2), and FNS-CD-X2CAMF-ADC(3) methods with experimental values for Ar, Kr, Xe, and Rn in the dyall.av4z basis set.
Atom State CCSD(3h1p)Mukhopadhyay et al. (2025b) ADC(2) ADC(3) Expt.Kramida et al. (2024)
Ar $^{3}P_{2}$ 0.24 -2.62 0.07 43.39
$^{3}P_{1}$ 0.24 -2.63 0.08 43.53
$^{3}P_{0}$ 0.25 -2.63 0.08 43.58
$^{1}D_{2}$ 0.27 -2.77 0.14 45.13
$^{1}S_{0}$ 0.35 -2.61 0.23 47.51
Kr $^{3}P_{2}$ 0.14 -2.14 -0.02 38.36
$^{3}P_{1}$ 0.16 -2.20 0.00 38.92
$^{3}P_{0}$ 0.18 -2.20 0.00 39.02
$^{1}D_{2}$ 0.18 -2.29 0.04 40.18
$^{1}S_{0}$ 0.23 -2.21 0.11 42.46
Xe $^{3}P_{2}$ 0.06 -1.73 -0.12 33.11
$^{3}P_{1}$ 0.06 -1.86 -0.12 34.32
$^{3}P_{0}$ 0.10 -1.74 0.09 34.11
$^{1}D_{2}$ 0.10 -1.89 -0.06 35.23
$^{1}S_{0}$ 0.14 -1.87 0.00 37.58
Rn $^{3}P_{2}$ 0.02 -1.39 -0.43 29.74
MAE 0.17 2.17 0.10
MAD 0.35 2.77 0.43
STD 0.09 0.39 0.14
RMSD 0.19 2.21 0.14
To understand the effect of basis sets on the calculated DIP values, we utilize the dyall.avxz (x = 2, 3, 4) series,Dyall (1998, 2002, 2006, 2016) which provides a systematically improvable basis set hierarchy. Additionally, to examine the role of diffuse functions, we consider both singly and doubly augmented versions of the dyall.av4z basis set. The results, presented in Table 2, illustrate the dependency of DIP values in different basis sets for $X\>^{3}\Sigma^{-}$, $a\>^{1}\Delta$, $b\>^{1}\Sigma^{+}$, and $c\>^{1}\Sigma^{-}$ states of the $\mathrm{(Cl_{2})}^{2+}$molecule. It is evident that increasing the basis set quality from double-zeta to triple-zeta leads to an average increase of approximately 0.2 eV in the DIP values for the
Table 4: Comparison of error in DIP values (in eV) for Cl2, Br2, HBr, and HI molecules using 4c-FNS-CCSD($3h1p$), FNS-CD-X2CAMF-ADC(2), and FNS-CD-X2CAMF-ADC(3) methods with experimental values in dyall.av4z basis set.
Molecule State CCSD(3h1p)Mukhopadhyay et al. (2025b) ADC(2) ADC(3) Expt.McConkey et al. (1994); Fleig et al. (2008); Eland (2003b); Yencha et al. (2004)
Cl2 $X^{3}\Sigma^{-}$ 0.47 -1.97 0.14 31.13
$a^{1}\Delta$ 0.35 -1.93 0.03 31.74
$b^{1}\Sigma^{+}$ 0.47 -1.75 0.04 32.12
$c^{1}\Sigma^{-}$ 0.54 -1.14 0.20 32.97
Br2 A $0_{g}$ 0.14 -1.74 0.00 28.39
A $1_{g}$ 0.13 -1.78 0.02 28.53
A $2_{g}$ 0.27 -1.58 0.04 28.91
A $0_{g}$ 0.16 -1.45 0.04 29.38
HBr $X^{3}\Sigma^{-}$ 0.38 -1.93 0.19 32.62
$a^{1}\Delta$ 0.43 -2.02 0.26 33.95
$b^{1}\Sigma^{+}$ 0.48 -1.95 0.34 35.19
HI $X^{3}\Sigma_{0}^{-}$ 0.11 -1.64 -0.11 29.15
$A^{3}\Sigma_{1}^{-}$ 0.12 -1.69 -0.11 29.37
$a^{1}\Delta$ 0.15 -1.71 -0.03 30.39
$b^{1}\Sigma^{+}$ 0.23 -1.64 0.03 31.64
MAE 0.30 1.73 0.11
MAD 0.54 2.02 0.34
STD 0.15 0.22 0.12
RMSD 0.33 1.74 0.14
four considered states. A further improvement from triple-zeta to quadruple-zeta results in a slightly smaller average increase of 0.16 eV. Unfortunately, quintuple-zeta dyall basis sets are not yet available for Cl.Reitsma et al. (2025) Interestingly, the addition of diffuse functions appears to have minimal impact: the singly augmented dyall.av4z basis set yields no noticeable change in DIP values, while the doubly augmented version results in a modest increase of only
Table 5: Comparison of DIP values (in eV) of Br2 with previous theoretical and experimental results.11footnotetext: Calculated at dyall.av3z basis. Taken from Pathak et al.Pathak et al. (2020)22footnotetext: Calculated at dyall.av4z basis using 4c-FNS approximation. Taken from Mukhopadhyay et al.Mukhopadhyay et al. (2025b)33footnotetext: FNS-CD-X2CAMF approximation is used.
State CCSD(3h1p)a CCSD(3h1p)b MRCIFleig et al. (2008) ADC(3)c Expt.Fleig et al. (2008)
A $0_{g}$ 28.53 28.39 28.39 28.39
A $1_{g}$ 28.47 28.66 28.54 28.55 28.53
A $2_{g}$ 29.04 29.18 29.01 28.95 28.91
A $0_{g}$ 29.52 29.54 29.45 29.42 29.38
B $0_{u}$ 29.88 29.78 29.72
B $3_{u}$ 29.91 29.81 29.77
B $2_{u}$ 29.79 30.26 30.16 30.12 30.30
B $1_{u}$ 30.33 30.24 30.19
B $0_{u}$ 30.60 30.50 30.45
B $1_{u}$ 30.62 30.52 30.48
0.01 eV per state. Thus, all further computations are performed using the dyall.av4z basis set, without adding any diffuse functions.
@@ IV.4 Comparison of DIP-ADC results with available theoretical and experimental data
To benchmark the performance of DIP-ADC(2), and DIP-ADC(3) methods, we have computed the DIPs of Ar, Kr, Xe, and Rn, for which corresponding experimental dataKramida et al. (2024) are available. The results are presented in Table 3, alongside 4c-FNS-DIP-EOM-CCSD($3h1p$) values and various statistical error metrics, with experimental data taken as the reference. For simplicity, we will hereafter denote DIP-EOM-CCSD($3h1p$) as CCSD($3h1p$) throughout this work. The statistical parameters reported include the mean absolute error (MAE), maximum absolute deviation (MAD), standard deviation (STD), and root mean square deviation (RMSD). It can be clearly seen that DIP-ADC(2) consistently underestimates the double ionization potentials for all the noble gas states studied, with a MAE of 2.17 eV. The largest deviation, 2.77 eV, is observed for the $^{1}D_{2}$ state of the Ar atom. This poor performance is not unexpected, as double ionization leads to a significant change in the reference ground state wave function, which results in a significant orbital relaxation effect in the doubly ionized states. A more precise treatment of the $3h1p$ excitations, achieved by incorporating higher-order terms in the ADC secular matrix, is essential for accurately capturing these effects. Since ADC(2) only includes zeroth-order terms in the 3h1p3h1p-$3h1p$ block, it fails to account for the necessary relaxation, resulting in large errors. The first level of improvement is seen with the ADC(2)-x method, which includes first-order corrections in the 3h1p3h1p-$3h1p$ block in an ad hoc manner. This lowers the MAE from 2.17 eV, observed in strict ADC(2), to 0.58 eV in ADC(2)-x (see Table S3 of the Supplementary Material). Among all the methods tested, DIP-ADC(3) shows the best agreement with experiment, with an MAE of only 0.1 eV and a MAD of 0.43 eV. In comparison, the 4c-FNS-CCSD($3h1p$) method yields a MAE of 0.17 eV and a MAD of 0.35 eV.
Table 4 presents the DIP values for the linear diatomic molecules-Cl2, Br2, HBr, and HI. In the ADC(2) method, a large MAE of 1.73 eV and a MAD of 2.02 eV are observed. In contrast, the significantly lower MAE of 0.30 eV obtained from the FNS-CCSD(3h1p) method highlights the importance of incorporating higher-order treatment of $3h1p$ excitations. This is further supported by the DIP-ADC(3) results, which demonstrate excellent performance with a MAE of just 0.11 eV and an RMSD of 0.14 eV across the 15 states examined in this study. The DIP-ADC(2)-x method yields a significantly lower MAE of 0.55 eV compared to the 1.73 eV observed with the strict ADC(2) method (See Table S4 of the Supplementary Material).
@@
IV.5 Different levels of relativistic hamiltonian treatment
Table 6: Comparison of FNS-CD-X2CAMF-DIP-ADC(3)/dyall.av4z results (in eV) for Ar, Kr, Xe, and Rn atoms across non-relativistic and various relativistic Hamiltonians.
Atom State NR DC DCG DCB Expt.Kramida et al. (2024)
Ar $^{3}P_{2}$ 43.56 43.46 43.45 43.45 43.39
$^{3}P_{1}$ 43.56 43.60 43.59 43.59 43.53
$^{3}P_{0}$ 43.56 43.66 43.65 43.65 43.58
$^{1}D_{2}$ 45.28 45.26 45.25 45.25 45.13
$^{1}S_{0}$ 47.75 47.74 47.73 47.73 47.51
Kr $^{3}P_{2}$ 38.65 38.34 38.33 38.33 38.36
$^{3}P_{1}$ 38.65 38.92 38.89 38.90 38.92
$^{3}P_{0}$ 38.65 39.02 38.99 39.00 39.02
$^{1}D_{2}$ 40.12 40.22 40.19 40.20 40.18
$^{1}S_{0}$ 42.37 42.58 42.55 42.55 42.46
Xe $^{3}P_{2}$ 33.60 32.98 32.97 32.97 33.11
$^{3}P_{1}$ 33.60 34.20 34.17 34.17 34.32
$^{3}P_{0}$ 33.60 34.02 34.00 34.00 34.11
$^{1}D_{2}$ 34.78 35.16 35.13 34.14 35.23
$^{1}S_{0}$ 36.70 37.58 37.54 37.54 37.58
Rn $^{3}P_{2}$ 31.14 29.31 29.29 29.30 29.74
MAE 0.38 0.10 0.11 0.10
MAD 1.40 0.43 0.45 0.44
STD 0.52 0.14 0.15 0.14
RMSD 0.52 0.14 0.15 0.15
Table 7: Comparison of FNS-CD-X2CAMF-DIP-ADC(3)/dyall.av4z results (in eV) for Cl2, Br2, HBr, and HI molecules across non-relativistic and various relativistic Hamiltonians.
Molecule State NR DC DCG DCB Expt.McConkey et al. (1994); Fleig et al. (2008); Eland (2003b); Yencha et al. (2004)
Cl2 $X^{3}\Sigma^{-}$ 31.32 31.27 31.27 31.27 31.13
$a^{1}\Delta$ 31.81 31.77 31.77 31.77 31.74
$b^{1}\Sigma^{+}$ 32.18 32.16 32.15 32.15 32.12
$c^{1}\Sigma^{-}$ 33.24 33.17 33.17 33.17 32.97
Br2 A $0_{g}$ 28.60 28.39 28.38 28.38 28.39
A $1_{g}$ 28.60 28.55 28.53 28.54 28.53
A $2_{g}$ 29.01 28.95 28.94 28.94 28.91
A $0_{g}$ 29.33 29.42 29.40 29.41 29.38
HBr $X^{3}\Sigma^{-}$ 32.89 32.81 32.80 32.80 32.62
$a^{1}\Delta$ 34.24 34.21 34.20 34.20 33.95
$b^{1}\Sigma^{+}$ 35.49 35.53 35.51 35.51 35.19
HI $X^{3}\Sigma_{0}^{-}$ 29.26 29.04 29.03 29.03 29.15
$A^{3}\Sigma_{1}^{-}$ 29.26 29.27 29.25 29.25 29.37
$a^{1}\Delta$ 30.37 30.36 30.34 30.34 30.39
$b^{1}\Sigma^{+}$ 31.47 31.67 31.65 31.65 31.64
MAE 0.15 0.11 0.10 0.10
MAD 0.30 0.34 0.32 0.32
STD 0.14 0.12 0.12 0.12
RMSD 0.18 0.14 0.14 0.14